OpenStudy (anonymous):

Bill has 5 times as many baseball cards as Fred. Time has 3 fewer than 4 times as many baseball cards as Fred. In all they have 167 baseball cards. How many does each have? Bill has 5 times as many baseball cards as Fred. Time has 3 fewer than 4 times as many baseball cards as Fred. In all they have 167 baseball cards. How many does each have? @Mathematics

OpenStudy (anonymous):

let x = {number of cards Fred has} then {number of cards Time has}=4x-3 and {number of cards Bill has}=5x {Fred}+{Time}+{Bill}=167 x+4x-3+5x=167 10x=170 x=17 {Fred}=17 {Time}=4(17)-3=65 {Bill}=5(17)=85 Check: 17+65+85=167